OpenStudy (anonymous):

y=2sqrtx-x find domain, horizontal &vertical asymptotes, symmetry, first derivative

OpenStudy (anonymous):

\[y=2\sqrt{x}-x\]

OpenStudy (anonymous):

helppppppppp plzzzz

OpenStudy (anonymous):

\[2 \sqrt {x-x}\] is it like that?

OpenStudy (anonymous):

i wrote it above

OpenStudy (anonymous):

2\[2{\sqrt{x}}-x\]

OpenStudy (anonymous):

Domain: x >= 0

OpenStudy (anonymous):

ok i got that, no horizontal or vertical asymptotes right

OpenStudy (anonymous):

?

OpenStudy (anonymous):

no, no asymptotes as the domain can be 0 and any x greater than 0. For DERIVATIVE you are going to use the subtraction law and the index law and rewrite it as follows \[2x^\frac{ 1 }{ 2 }\]

OpenStudy (anonymous):

waittttttttttttttttttt

OpenStudy (anonymous):

can i tell u wat i got for the derivative

OpenStudy (anonymous):

\[\frac{ d}{ dx }2x^\frac{ 1 }{ 2 } - \frac{ d }{ dx }2\]

OpenStudy (anonymous):

is your derivative \[x^\frac{- 1 }{ 2 }\]

OpenStudy (anonymous):

?

OpenStudy (anonymous):

quick i need to go.

OpenStudy (anonymous):

i got x^1/2 -1

OpenStudy (anonymous):

u forgot derivative of x

OpenStudy (anonymous):

its y=2√x -x

OpenStudy (anonymous):

my bad, i replaced x by 2 in my equation. Just change yours to have a NEGATIVE in the power. So NEGATIVE 1/2.

OpenStudy (anonymous):

remember derivative n.a^n-1

OpenStudy (anonymous):

ok i got it

OpenStudy (anonymous):

it must be x^-1/2 - 1

OpenStudy (anonymous):

cheers